OpenStudy (xximhisgirlxx):

If BC bisects ∠ABD, and ∠ABC = 5x + 11, ∠CBD = 10x − 4, then find x 5 3 1 2

OpenStudy (ayyookyndall):

B

OpenStudy (xximhisgirlxx):

How did you get B?

OpenStudy (ayyookyndall):

5x+11 = 10x-4 -5x+11=-4 -5x=-15 x=3

OpenStudy (campbell_st):

well bisecting means that the line BC cuts the angle ABD into 2 equal parts. so the angles are equal in size then you need to solve 5x + 11= 10x - 4 hope it helps